I am thinking of a number. If you multily my number by 2 and then add 3, you will get 55. What is my number?
Answer by checkley75(3666)

2N+3=55
2N=55-3
2N=52
N=52/2
N=26 ANSWER FOR YOUR NUMBER.
PROOF
2*26+3=55
52+3=55
55=55